Two countries, Great Britain and the United States, produce just one good: beef. Suppose the price of beef in the United States is $2.80 per pound and in Britain it is £3.70 per pound.

  1. According to PPP theory, what should the dollar/pound spot exchange rate be?

  2. Suppose the price of beef is expected to rise to $3.10 in the United States and to £4.65 in Britain. What should the one-year forward dollar/pound exchange rate be?

  3. Given your answers to parts a and b, and given that the current interest rate in the United States is 10 percent, what would you expect the current interest rate to be in Britain?
Answer #1

Price of Beef in US P1 = $2.80

Price of Beef in Britain P2 = $3.70

  1. According to PPP theory i.e. comparision of different countries currencies given through a basket of goods approach.

Therefore,

Exchange rate of curreny USA to currency Britain is calculated as,

S= P1 /P2

= 2.8/ 3.7

= 0.7567

Hence, Spot exchange rate is 0.7567

  1. Revised expected prices are

P1’ = 3.1

P2’ = 4.65

Exchange rate given by,

F = P1’/ P2’

= 3.1/ 4.65

= 0.667

Hence, one-year forward dollar/pound exchange rate is 0.667

  1. According to interest rate parity,

Let, ic and ib be the interest rate for US & Britain resp. The interest rate in US is given as 10%. i.e. ic= 10%= 0.1

And F0 and S0 forward and Spot exchange rate of US.

0.667 = 0.7567 * (1+.1)/(1+ib)

On solving, we get

ib = 24.86 %

Hence, interest rate of Britain is 24.86%

If you like the answer please give an up-vote, this will be quite encouraging for me, thank you.

If you have further suggestions in answer, please write below on the comment section, I will certainly look into it.